OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (anonymous):

i got 3x+y=16 is that correct???

terenzreignz (terenzreignz):

No. To test out, it has to work with both points. Try P where x = 6 and y = 2 3(6) + 2 = 16 ? I think not ;)

OpenStudy (anonymous):

ohh nvm its 3x+y=20